### v2

# Multi-Modal Framing Analysis of News

Abstract
Automated frame analysis of political communication is a popular task in computational social science that is used to study how authors select aspects of a topic to frame its reception. So far, such studies have been narrow, in that they use a fixed set of pre-defined frames and focus only on the text, ignoring the visual contexts in which those texts appear. Especially for framing in the news, this leaves out valuable information about editorial choices, which include not just the written article but also accompanying photographs. To overcome such limitations, we present a method for conducting multi-modal, multi-label framing analysis at scale using large (vision-)language models. Grounding our work in framing theory, we extract latent meaning embedded in images used to convey a certain point and contrast that to the text by comparing the respective frames used. We also identify highly partisan framing of topics with issue-specific frame analysis found in prior qualitative work. We demonstrate a method for doing scalable integrative framing analysis of both text and image in news, providing a more complete picture for understanding media bias.
Multi-Modal Framing Analysis of News
Arnav Arorathanks:   denotes equal contribution      Srishti Yadav11footnotemark: 1 Maria Antoniak      Serge Belongie      Isabelle Augenstein University of Copenhagen {aar, srya, maan, s.belongie, augenstein}@di.ku.dk
1 Introduction
Frames are conceptual tools that both communicators and audiences use to interpret and categorize issues and events Gitlin (1980); Eko (1999); Pan and Kosicki (1993); Reese et al. (2001). By highlighting specific elements of a topic and minimizing others, communicators frame messages in ways they believe will resonate with audiences Goffman (1974) and can shape the way the topic is perceived by readers or viewers (Schudson, 2003). In the field of journalism, framing is a core narrative device by which news consumption is framed within an interpretive perspective (Card et al., 2015). Most prior work on computational frame analysis has focused on linguistic structure and content analyses via text elements (Ali and Hassan, 2022). However, framing is not solely textual; visual elements also play a crucial role in conveying implicit and explicit messages (Messaris and Abraham, 2001).
Figure 1: News can be intentionally framed to affect reader perception. Editorial choices decide what is communicated through the words and the images. Our approach systematically detects this framing.
News media often employ images alongside text to reinforce or contrast the intended frame, leveraging the affective and cognitive impact of visuals Cope et al. (2005); Keib et al. (2018); Grabe and Bucy (2009); Geise et al. (2025). An example can be seen in Figure 1. The image depicts protesters with signs with police presence around, using the public opinion and security framing. The article text, on the other hand, talks about the quality of life of migrants, economic implications, crime, and policy framing. Such differences in portrayal can affect readers’ perception of the crisis and their induced emotions. Further, while communication in text is more explicit through linguistic framing, images encode frames in more subtle and implicit ways, requiring sophisticated interpretation models to capture their meaning Aiello and Parry (2019). As such, when conducting media analysis, the communicated framing across both the image and the text should be considered.
While visual framing analysis has been explored in communication science and journalism studies (Wessler et al., 2016; Powell et al., 2017), existing studies doing computational framing analysis have largely ignored this crucial aspect (Ali and Hassan, 2022). Further, they have primarily focused on a fixed set of labels for framing analysis, performing prediction in a multi-class setting. This substantially limits the information one can derive from predictions as an article can convey several frames (Figure 1) and fine-grained analysis of framing within a topic necessitates frames specific to a particular issue. Large vision and language models are particularly well suited for conducting this task at scale, considering the semantic understanding embedded into them through large scale pre-training. Our contributions include the following.
  • We present the first computational study of multimodal framing in the news, outlining a methodology using LLMs and VLMs.
  • We provide a large-scale dataset  [^1] https://huggingface.co/datasets/copenlu/mm-framing of 500k U.S.-based news articles for framing analysis, automatically labeled with generic and issue-specific frames, validated through a smaller set of human annotations.
  • We conduct a thorough analysis of frames communicated in the image vs the text of articles, both at a corpus-level and at a fine-grained level for articles about immigration.
  • We discover meaningful differences in how frames are used in images and texts, across political leanings and topics.
2 News Framing
Framing of news articles has been studied widely in communication studies. There are several definitions of framing and scholars often disagree on the method for extraction and its role in public communication, leading to it being termed a “fractured paradigm” (Entman, 1993). The definition from Entman for framing at large, however, is the most widely accepted one. He defined framing as “making some aspects of reality more salient in a text in order to promote a particular problem definition, causal interpretation, moral evaluation, and/or treatment recommendation for the item described”. Narrowing the framework to news,  De Vreese (2005) lays out a typology of news framing. He states that there are generic news frames and issue-specific news frames. Generic news frames “transcend thematic limitations and can be identified in relation to different topics, some even over time and in different cultural contexts”. They are particularly useful for uncovering broad patterns within or across countries. Issue-specific frames allow for richer, more fine-grained analysis of various aspects highlighted within a particular issue.
2.1 Framing through visuals
Photographs are an important vehicle of framing as a reader may process textual and visual messages differently. Readers may focus on photographs without also reading an accompanying story Miller (1975) or might select which news story to read depending on the image thumbnail. Images are potent framing tools because they evoke immediate emotional responses, provide contextual cues, and sometimes contradict textual narratives Geise and Xu (2024). For visual data, however, automation is more challenging because computational image analysis often struggles with connotative and symbolic elements that are readily discernible to human annotators Rodriguez and Dimitrova (2011). While some recent advancements in machine learning and computer vision (e.g., object detection, facial recognition, sentiment analysis on images) offer promising avenues for automated visual framing analysis, the complexity of symbolic and ideological meanings typically requires human interpretation Coleman (2010).
Understanding multimodal framing is important for several reasons: it allows for a more granular examination of media bias, as textual analysis alone may miss the ideological and emotional undertones of visual elements Wessler et al. (2016); Geise et al. (2025). It improves fact-checking and detection of misleading content by identifying inconsistencies between textual claims and their supporting images. It also finds applications in political communication, journalism studies, and public policy analysis through tracing changes in framing over time and across outlets Baumgartner et al. (2008).
2.2 Integrative Framing Analysis
Integrative framing analysis is when both images and text are observed separately, but the results are integrated to form a more complete picture of framing analysis (Dan, 2017). Although there is broad agreement that visual and verbal elements should be studied side by side (Coleman, 2010; Rose, 2022), relatively few studies have effectively combined the two. Researchers often focus on textual framing Matthes and Kohring (2008) or visual framing Rodriguez and Dimitrova (2011) as separate domains, even though media messages typically blend both. Geise et al. (2013) and Coleman (2010) emphasize that analyzing visuals in isolation can obscure how words shape viewers’ interpretations—and vice versa. As computational tools become increasingly sophisticated and as scholarly interest in visual framing continues to grow, integrative approaches are poised to become more common. By integrating these modalities, scholars can better observe whether text and visuals reinforce each other or present divergent frames.
2.3 Automated Framing Analysis
In computational studies, framing has been studied with the help of machine learning methods for content analysis at scale.  Ali and Hassan (2022); Vallejo et al. (2023); Otmakhova et al. (2024) survey these efforts towards on computational framing analysis, focusing on different aspects like methods, varying conceptualisations and inter-disciplinary connections of framings. The most widely used resource is one by  Card et al. (2015), who present the Media Frames Corpus (MFC), a dataset of US based news headlines annotated for generic frames. Most studies use it in a supervised, multi-class prediction setting with MFC frames for analysis of social networks (Mendelsohn et al., 2021) or discussion forums (Hartmann et al., 2019). There are also English-only and multi-lingual SemEval tasks on frame detection. (Piskorski et al., 2023a, b; Sajwani et al., 2024). Unsupervised approaches to framing analysis use framing lexica (Field et al., 2018), clustering (Burscher et al., 2016; Ajjour et al., 2019), and topic models (Nguyen, 2015). All of these approaches have focused on texts alone.
Figure 2: Distribution of data across the top 30 topics
3 Dataset
We crawl news articles along with corresponding images from 28 US-based news agencies. We extract data for a 12 month period between May 2023 and April 2024 using the newsplease library (Hamborg et al., 2017). Our selected news sources reflect the entire political spectrum, based on data from AllSides Media Bias, [^2] https://www.allsides.com/media-bias/ratings an organization that assigns a rating of Left, Lean Left, Center, Lean Right, or Right to each media outlet. Our list of sources along with corresponding political leanings are listed in Appendix A. We query the publicly available Common Crawl archives [^3] https://commoncrawl.org/news-crawl for the corresponding publishers and extracted each article’s text, headline, publication date, image_urls and other metadata in JSON format. Post scraping, we filter extremely short and long articles, images with logos and other noise. Further details are provided in Appendix B. Our final dataset includes  500K articles and corresponding images. We show the distribution of the data across the time period of data collection in Figure 3.
Figure 3: Distribution of data across the time-period of collection, broken down by political leaning.
4 Method
4.1 Model Annotation
We use Large Language Models (LLM) and Vision-Language Models (VLM) to label several aspects of the text and image from news articles, respectively. In Table 1, we list the different aspects extracted per modality from the news articles. We also generate explanations for the decisions along each prediction for reducing hallucination and facilitating downstream qualitative analysis. For extracting generic frame across both modalities, we use the framing dimensions outlined by Boydstun et al. (2014), as provided in the Media Frames Corpus (Card et al., 2015), which includes 15 generic frames appropriate for analysis of US news. These frames are Economic, Capacity & Resources, Morality. Fairness & Equality, Legality, Constitutionality & Jurisprudence, Policy Prescription & Evaluation, Crime & Punishment, Security & Defense, Health & Safety, Quality of Life, Cultural Identity, Public Opinion, Political, External Regulation & Reputation and Other. The descriptions of each of the frames are provided in Appendix C. Differing from most previous studies on automated framing analysis, which assign a single frame to an article, we conduct the frame extraction in a multi-label setting, i.e., an article or an image can have more than one frame. This setting is much closer to the setup scholars use when conducting qualitative studies on framing (Dan, 2017). For the main subject of the articles, the models were instructed to extract entities central to the text or the image, if there is one.
Modality Aspect Description
Text Topic Broad topic of the article
Main subject Subject of the text
Generic Frame One or more of the 15 generic media frames
Issue Frame Inductive, open-ended framing of a topic
Image Caption Caption of the image
Main Subject Subject in the image
Generic Frame One or more of the 15 generic media frames
Table 1: List of aspects extracted per modality.
4.2 Models and Prompts
For text annotations, we use Mistral-7B. For images, we use Pixtral-12B (multi-modal model). We use the vLLM library for optimized memory usage and high-throughput inference, allowing us to conduct analysis on the entire dataset in a few days, demonstrating our approach’s scalability. See Appendix D for full modelling details.
We carefully craft prompts for the extraction of each aspect from the article. For extraction of frames expressed in the articles and images, we experiment with including definitions of framing and descriptions of each frame at several granularities. We benchmark these different prompts on a validation set. We also iteratively improve the prompt by looking at the category of errors and including those in the prompt. For instance, the Pixtral model has a tendency to predict the economic frame every time an entity in professional attire appears in the image, associating professional attire with being wealthy. We instruct the model to avoid such reliance. Finally, we report performance on a held-out test set. For issue-specific frames, we prompt the model by providing a definition of framing and some examples of issue-specific frames across topics. The task was then to analyse the article and generate an issue-specific frame with respect to the topic of the article. The full list of prompts for the extraction of each aspect is provided in Appendix  1 and  2.
Figure 4: Frequency of predicted generic frames across all articles for texts and images.
4.3 Model Evaluation
Text Framing
To assess the quality of output frames from our approach, and compare against prior work, we evaluate our text frame extraction model on the existing large scale benchmark Media Frames Corpus (Card et al., 2015). This corpus provides frame labels by each annotator for over 32,000 articles from US news on topics like immigration, smoking, and same-sex marriage. We take the union of all frame labels assigned by the annotators and assign the up to the top 3 most frequent frames assigned by the annotators as the label for an article. We run our text framing analysis model on this dataset, providing the article of the text while allowing the model to generate multiple frames per article. To assess model performance, we calculate the intersection between the sets of model-annotated frames and human-annotated frames for each article in the dataset. 95.7% of the articles had a non-zero intersection, with at least one overlapping frame label, demonstrating that the model outputs accurate frames in most cases. Our model received a micro averaged F1 score of 0.5, with an averaged precision of 0.42 and averaged recall of 0.62. We include the scores per frame label in Appendix F. For prediction across 15 labels and a task as subjective as framing, we believe the model performs quite well. We further did qualitative assessment of model outputs by going through the model explanations for the predicted frames and found the model to be competent.
Image Framing
For images, there is no existing benchmark with frame labels. Two of the authors of this study manually annotated 600 images for generic frames across a stratified (time-period and topics) sample of the dataset. We modify generic frame annotation guidelines released by Mendelsohn et al. (2021), adapting them to the image annotation setting. We set up a multi-label classification platform (as seen in Figure 12) where annotators are allowed to select one or more frames from the list given an image. Framing analysis is a highly subjective task, as is well established by prior work in communication studies as well as NLP. For images, the subjectivity increases substantially, given the limited amount of context available and requiring cultural and conceptual familiarity more so than needed for textual framing analysis (Geise and Baden, 2015). We then calculate the proportional overlap in frames between annotators, counting a non-zero intersection of frame annotations as 1 and zero intersections as 0. The proportional overlap between the annotators is 85%, demonstrating the high percentage of instances with overlapping frame labels. Using this dataset, we create an evaluation set for our vision model and get predictions from our pixtral model. Since ’Sports’ news are mostly labelled as ’cultural identity’, we remove them (8%of them) for our calculation. We take the union of the frame predictions for the two annotators and calculate the intersection with the model predictions for a particular image. The proportion of non-zero intersection instances between the image and human annotations is 84.2%, i.e., at least one correctly predicted frame most of the time.
Topics
To evaluate the topics output by the model, we use the 19 most frequent topics for our analysis, discarding one of the top 20 topics (Media) because most of its included documents were publication boilerplate rather than article texts. Two of the authors hand-annotate a set of 190 articles (10 per topic), marking whether the model’s topic prediction was acceptable. The overall accuracy for topics predicted by the model is 86% and 87% per annotator. We discard two topics (Business, Environment) with the lowest accuracy scores, leaving us with a final set of 17 topics. We list these, with examples and accuracies, in Appendix Table 7.
Figure 5: Comparison of generic frame prediction frequencies in images vs texts. The x-axis represents the subtracted rank of predictions for that frame between the two modalities: positive scores indicate that the frame was more often predicted in texts, negative scores indicate that the frame was more often predicted in images.
5 Multi-Modal Framing Analysis
5.1 Generic Framing Analysis
We first analyze the generic frames predictions across the dataset [^4] For framing analysis, we perform additional filtering, excluding articles with “None” or other non-valid frame predictions. We further removed articles less than 100 words and articles on the topics ”sports” and ”media”, as we observed while conducting qualitative analysis that articles on sports only used the Cultural Identity frame in the images and articles with the media topic were of generally low quality with several pages only including text from a paywall, possibly skewing the results. This resulted in a reduced set of  154k articles in  Figure 4. Overall, we can see that there are many more predicted frames for text compared to the images (a mean of 3.6 predicted frames per text and 1.3 predicted frames per image). This is intuitive (and can be seen in the  Figure 1) since the text of the article can more easily express several distinct frames. Looking at their distributions, we can see that in the text of the article, quality of life, crime, and policy appear more frequently, while morality and external regulation & reputation are more rare. In terms of image frame distribution, external regulation & reputation has very low frequency. Contrasting the two modalities, economic and policy frames more frequently appear in text, while capacity & resources and culture are more prevalent in images.
We also compare frames by modality across topics, as shown in Figure 5. We observe substantial differences in how topics are framed across images and texts. When covering war, news outlets focus more on the external regulation & reputation and crime framing in the text, but the images that are used convey public opinion and capacity & resources framing, depicting people giving speeches and of military equipment. For articles on the topic of economy, there is more focus on fairness and security & defense spending in texts, while the images depict the health and safety, culture, and political frames.
Figure 6: Pointwise mutual information between text and image frames across the dataset. Some frames are used consistently in texts and images for the same article (dark cells in the diagonal), other frames differ widely.
Frame Co-Occurrence
To understand what is highlighted in the text of the article compared to the image, we plot the pointwise mutual information (PMI) of image frames and text frames across the entire dataset in Figure 6. We can see the presence of a diagonal, demonstrating that there is often alignment between frames across modalities. However, there are many deviations, some of which are intuitive. For example, depictions of quality of life in the images when writing about cultural topics and vice versa is quite common. Quality of life framing in the images is also used when the text is using an economic frame. Political framing in the text is associated with policy and public opinion framing, and legal framing in text is associated with political framing in the images. To get a finer-grained understanding of co-occurrence of frames across the modalities, we analyse the percentage of co-occurrences per topic, as shown in Figure 7. For articles about crime, when using the criminal framing in images, the texts also tend to highlight the security, quality of life, and the legal frames. For war, the images consistently highlight the security & defense framing, even when the article text is highlighting the policy or legality frames.
Figure 7: Frequency of frame co-occurrence between text and image frames across four selected topics.
Lexical Comparison
crime quality of life
$z$-score bigram $z$-score bigram
Table 2: The bigrams most associated with the images (higher $z$-scores) and texts (lower $z$-scores) for the crime and quality of life frames.
Our results above indicate that frames are used in images and texts in different ways, but how those uses differ is unclear. Why might the same article include a crime frame for its text and a political frame for its image? To explore this question, we perform a lexical analysis of the words used in articles whose image or text used the same frame. We use the “Fightin’ Words” algorithm from Monroe et al. (2008), a comparison metric which takes into account disproportionate numbers of samples as well as rare words, to find the bigrams from the article texts most associated with a single frame for images vs texts. Table 5.1 shows the bigrams sorted by their $z$-scores (prior=$0.01$, frequency5absent5$\geq$ 5≥ 5) for two frames. Qualitatively, we observe that words associated more with image frames tend to be concrete (“ice cream”) and associated with a single meaning more easily recognized to the predicted frame (“police department”). Prior work suggests that more tightly clustered and recognizable images are associated with more concrete topics (Hessel et al., 2018). See the Appendix for a full list of the frames and sorted bigrams.
Figure 8: Comparison of text and image frame distributions across political leanings for five topics.
Political Leaning
We also analyzed the correlation between political leaning and text/image frames across topics. For each topic, we compute how often each frame appears for each political leaning, combining left-lean and left as well as right-lean and right (see §3). Figure 8 shows the proportional frequency of frames per leaning, allowing us to compare which frames dominate among different leanings for a specific topic. We observe that for topics like crime and war, the images do not portray many frames other than frames like crime and security, unlike texts where we see additional frames like legality, policy, regulation and some references to the political frame across all political spectrum. On the other hand, topics like immigration have more variation. Right-leaning news agencies focus more on legality, policy, and security frames in text and on public opinion and security frames in images. Similarly for politics, apart from the political frame, images unlike texts do not often feature the crime frame.
Figure 9: Examples of generic frame prediction in images vs texts about immigration across political leanings.
5.2 Case Study: Immigration
So far, we demonstrated the differences in generic framing across across modalities, topics, and political leaning across an entire corpus of articles. But our approach also allows for a more fine-grained analysis, and we demonstrate this in a case study focused on one topic: immigration.
In Figure 5, we show that articles about immigration use the crime and punishment, external regulation & reputation, and economic framing much more frequently than the images. These articles often focus on deals with other countries, their contribution to the economy, the cost of their deportation, and/or the crimes that they commit. On the other hand, images tend to use the capacity & resources framing, showing migrants in camps, or the public opinion framing, showing people giving speeches. There are also differences across the political spectrum, as can be seen in Figure 8, with the right focusing much more on the security & defense framing compared to sources from the left or center across both modalities. To highlight this further, we show the finer-grained issue specific frames per political leaning in Figure 10. We see clear differences and fine-grained signals about which specific framing publishers leaning to different sides of the political spectrum use. The left and center tend to highlight the humanitarian crisis framing much more, the gap is smaller for right leaning publishers, while on the right, the focus is more on immigrants being an economic burden or a security threat.
Figure 10: Percentage of issue frames used by news publishers in the article text across the political spectrum in their reporting on immigration.
6 Discussion
Being exposed to selective information can significantly bias our world view. For news, it can lead to problems at an individual level such as miscommunication; at a societal level, it can lead to misinformation, political polarisation, or avoidance of news (Lecheler, 2018; Iyengar, 2017). This problem is aggravated considering news consumers tend to receive their news from social media, further creating “curation bubbles” (Green et al., 2025), where the dynamics of the networks on social media further contribute to curation, and in effect, exposure to selective partisan information. Thus, it is important to conduct analyses of news framing at scale to understand its effects (Lecheler, 2018).
Restricting computational analyses of framing to texts only or single labels alone results in missing information that could reshape one’s interpretation of an article. An article can use several frames within the text itself. Further, analysing images is important because it allows for a more complete analysis of media bias, as textual analysis alone may miss the ideological and emotional undertones of visual elements Wessler et al. (2016); Geise et al. (2025). While analyzing images for frames is difficult, as the task is subjective, we find that it is possible reach a meaningful agreement between trained annotators and detect frames with reasonable accuracy across large and diverse datasets.
Our analysis of multimodal framing in the news reveals that more frames are predicted for texts compared to images, that there are distinct lexical patterns associated with text frames vs. image frames, and that frames can be used for detailed topic analysis, including both at a corpus level and at a more fine-grained level by focusing on a single issue. Finally, we show how different news publishers from different ends of the political spectrum use different issue-specific frames to build a narrative, portraying different sides of a story to their readers. The differences we uncover between the use of text and image frames emphasize that it is crucial to take images into account when analyzing framing in the news.
7 Conclusion
Framing analysis is an important task, with implications for several fields in the social sciences and for understanding society at large. Through our work, we provide a dataset and leverage large vision-language and language models to conduct a more integrative framing analysis, which we hope is useful for future studies in framing analysis in computational and social sciences. This allows for more efficient and replicable analyses of large-scale media corpora. We also urge practitioners working on these topics to look at framing at a broader level and conduct multi-faceted analyses, taking into account both generic and issue-specific frames, visual framing, and portrayal of subjects in their data. Future research should explore human-in-the-loop approaches to ensure both efficiency and interpretive depth. As digital media continues to expand, automated approaches will be crucial for capturing and analyzing the evolving landscape of visual and textual framing across the news ecosystem.
8 Acknowledgments
This work is partially funded by a DFF Sapere Aude research leader grant under grant agreement No 0171-00034B, as well as by the Pioneer Centre for AI, DNRF grant number P1.
9 Limitations
Despite the advancements in automating multimodal framing analysis, our approach has certain limitations. Firstly, our method assumes that both textual and visual elements contribute meaningfully to framing, yet in some cases, framing might be driven predominantly by one modality. This is akin to a news reader focusing more on one modality while consuming news. The challenge of evaluating the dominance of one modality remain an open problem when text and image signals conflict or reinforce different interpretations. Secondly, the interpretability of automated framing predictions remains limited. While our approach can identify patterns aligned with theoretical framing literature, it does not fully replicate the depth of qualitative human analysis. Ensuring transparency in model decisions and improving explainability remains an important avenue for future research. Finally, our study focuses on a specific set of framing methodologies and datasets, meaning generalizability to other media contexts or platforms requires further investigation. The dynamic nature of framing, influenced by evolving cultural and political landscapes, suggests that models must be continuously updated to remain effective.
10 Ethical Considerations
The models used in this study have been trained on large-scale datasets that may encode existing societal biases, which can impact the fairness and accuracy of framing predictions. If left uninvestigated, these biases may reinforce dominant narratives while marginalizing alternative perspectives, particularly in politically sensitive or socially divisive topics. Though having a scalable and automated approach can help aid news organizations and journalism scholars, it must be kept in mind that the generated responses are based on probabilistic patterns, hence we should allow error margins for some misclassifications. News is dynamic in nature but models can become static in nature, hence any method for news framing necessitates continuous updates of these large models.
Appendix A Source selection
We provide the list of all sources along with their corresponding political leanings in Table 3.
Leaning News Domain
Left alternet.org, editor.cnn.com, democracynow.org, dailybeast.com, huffpost.com, theintercept.com, jacobin.com, motherjones.com, newyorker.com, slate.com, msnbc.com, vox.com
Left Leaning abcnews.com, apnews.com, theatlantic.com, bloomberg.com, cbsnews.com, insider.com, nbcnews.com, thenytimes.com, npr.com, politico.com, propublica.org, time.com, washingtonpost.com, yahoonews.com, usatoday.com, theguardian.com
Center axios.com, forbes.com, newsweek.com, reuters.com, realclearpolitics.com, thehill.com
Right Leaning thedispatch.com, theepochtimes.com, foxbusiness.com, ijr.com, nypost.com, thepostmillennial.com, washingtonexaminer.com, washingtontimes.com
Right theamericanconservative.com, theamericanspectator.com, breitbart.com, dailycaller.com, dailywire.com, foxnews.com, newsmax.com, oann.com, thefederalist.com
Table 3: List of news sources used for our dataset split by their political leaning
Label Precision Recall F1-score
cap&res 0.39 0.34 0.36
crime 0.50 0.87 0.63
culture 0.38 0.37 0.37
economic 0.43 0.69 0.53
fairness 0.17 0.74 0.28
health 0.48 0.48 0.48
legality 0.53 0.87 0.66
morality 0.30 0.63 0.41
policy 0.40 0.73 0.51
political 0.68 0.53 0.60
public_op 0.32 0.55 0.40
quality_life 0.28 0.36 0.31
regulation 0.26 0.48 0.34
security 0.30 0.45 0.36
micro avg 0.42 0.62 0.50
macro avg 0.39 0.58 0.45
weighted avg 0.45 0.62 0.51
samples avg 0.46 0.63 0.51
Table 4: Metrics per frame label (multi-label) for our text frame classifier on the MFC dataset
legality morality security regulation
$z$-score bigram $z$-score bigram $z$-score bigram $z$-score bigram
more associated with images
Table 5: The bigrams most associated the images (higher $z$-scores) and texts (lower $z$-scores) for all the frames except crime and quality of life (shown in Table 5.1).
Appendix B Data Filtering
The scraped data carried a lot of noise. There were single line or extremely long articles, images that only depicted logos of news websites, which are not useful for analysis. To remove these, we filtered the articles whose lengths below the 5th percentile and above the 95th percentile, removing the outliers. For images, we similarly removed the files for which the size was above the 95th percentile. There was no clear way to remove the images with logos. To approximate this, we remove all figures with a size of less than 5000 bytes. We also removed articles that were not in English, an information available via news-please library.
Appendix C Frame Descriptions
In Table 6, we provide the names and corresponding descriptions for the generic frames used in our work. There are minor differences in the frame names provided here and the ones used in our dataset. This was due to us relying on the version released by Card et al. (2015) with the Media Frames Corpus dataset, which builds on the same dimensions.
Appendix D Experimental Details
Mistral: We used Mistral-7B-Instruct-v0.3 available via hugging face  [^5] https://huggingface.co/mistralai/Mistral-7B-v0.3. We used vLLM for high-throughput and memory-efficient inference and set the parameters as temperature=0.2, max_tokens=4000, dtype=’half’ and max_model_len=8096.
Pixtral: We used Pixtral-12B-2409 available via hugging face  [^6] https://huggingface.co/mistralai/Pixtral-12B-2409. Similar to Mistral, we used vLLM and set the parameters as: temperature=0.2, max_tokens=1024, dtype="half" and max_model_len=7000. Before processing the images via pixtral, we also resized the images to 512 x 512 for computational efficiency.
For vision annotations, we ran pixtral on 8 Nvidia A100s which enabled us to finish the computation in 5 days. For text annotations, we ran mistral on a mix of Nvidia TitanRTX and A100s and finished the computation in similar time as vision annotations.
Frame Name Description
Economic The costs, benefits, or monetary/financial implications of the issue (to an individual, family, community or to the economy as a whole)
Capacity and resources The lack of or availability of physical, geographical, spatial, human, and financial resources, or the capacity of existing systems and resources to implement or carry out policy goals.
Morality Any perspective—or policy objective or action (including proposed action)—that is compelled by religious doctrine or interpretation, duty, honor, righteousness or any other sense of ethics or social responsibility.
Fairness and equality Equality or inequality with which laws, punishment, rewards, and resources are applied or distributed among individuals or groups. Also the balance between the rights or interests of one individual or group compared to another individual or group.
Constitutionality and jurisprudence The constraints imposed on or freedoms granted to individuals, government, and corporations via the Constitution, Bill of Rights and other amendments, or judicial interpretation. This deals specifically with the authority of government to regulate, and the authority of individuals/corporations to act independently of government.
Policy prescription and evaluation Particular policies proposed for addressing an identified problem, and figuring out if certain policies will work, or if existing policies are effective.
Law and order, crime and justice Specific policies in practice and their enforcement, incentives, and implications. Includes stories about enforcement and interpretation of laws by individuals and law enforcement, breaking laws, loopholes, fines, sentencing and punishment. Increases or reductions in crime.
Security and defence Security, threats to security, and protection of one’s person, family, in-group, nation, etc. Generally an action or a call to action that can be taken to protect the welfare of a person, group, nation sometimes from a not yet manifested threat.
Health and safety Healthcare access and effectiveness, illness, disease, sanitation, obesity, mental health effects, prevention of or perpetuation of gun violence, infrastructure and building safety.
Quality of life The effects of a policy on individuals’ wealth, mobility, access to resources, happiness, social structures, ease of day-to-day routines, quality of community life etc.
Cultural identity The social norms, trends, values and customs constituting culture(s), as they relate to a specific policy issue
Public opinion References to general social attitudes, polling and demographic information, as well as implied or actual consequences of diverging from or getting ahead of public opinion or polls.
Political Any political considerations surrounding an issue. Issue actions or efforts or stances that are political, such as partisan filibusters, lobbyist involvement, bipartisan efforts, deal-making and vote trading, appealing to one’s base, mentions of political manoeuvring. Explicit statements that a policy issue is good or bad for a particular political party.
External regulation and reputation The United States’ external relations with another nation; the external relations of one state with another; or relations between groups. This includes trade agreements and outcomes, comparisons of policy outcomes or desired policy outcomes.
Other Any frames that do not fit into the above categories.
Table 6: Frame name and descriptions for each frame used in the dataset
Appendix E Article Subject Portrayal
We additionally explore how individual entities in the articles are portrayed. For this analysis, we take the subset where the article and the image are portraying the same main entity. The entities and sentiment are extracted by prompting the LLM and VLM to identify the main subject in the article or the image, as shown in the prompt in 1. Figure 14 shows peculiar differences when contrasting the image portrayal from the portrayal in the articles. A general pattern that can be observed is that entities were portrayed more positively in the images, compared to the text, e.g., Rishi Sunak, Prince Harry, Elon Musk, Joe Biden, Benjamin Netanyahu. For some, there is a negative portrayal overall that is much more explicit in the text, e.g., Donald Trump, Hunter Biden, Vladimir Putin, Rudy Guliani.
Appendix F Frame Classifier Performance
We provide the per label performance of the framing classifier for text on the MFC dataset in Table 4.
Appendix G Additional Plots
Figure 11: Overall comparison of text and image frame distributions across political leanings
Figure 12: Image annotation UI used for annotating images frames by annotators
(a) images
(b) texts
Figure 13: Image and text frames cover different parts of the topic space. UMAP reduction (n_neighbors=200, cosine distance) of a 5k sample of the generated topic descriptions (TF-IDF weighted token vectors: max_features=5000, min_df=5, max_df=0.95) of the articles. Highlighted points represent the political frame.
Figure 14: Difference in sentiment of the main subject’s portrayal across text and images
SYS_PROMPT = f"""You are an intelligent and logical journalism scholar conducting analysis of images associated with news artciles."""
prompt_entity = """Your task is to identify the main subject or entity in the image that is central to the image. Entities are people or organizations that play a central role in the image. If there are multiple entities, choose the one that is central to the image. if there are no central or clearly identifiable entities, choose "None".
Additionally, analyse the image and output the sentiment with which the subject is portrayed in the image. The sentiment can be "positive", "negative" or "neutral". In case of no entity, output "None. Output only the json and no other text.
Format your output as a json entry as follows:
{"entity-name": "<name of the entity>", "sentiment": "<sentiment towards the entity>", "sentiment-reason": "<reasoning for the portrayed sentiment>"}
<image>
For the image above, provide the name of the entity, sentiment towards the entity, and reasoning for the chosen sentiment."""
FRAMES = f"""
Economic - costs, benefits, or other finance related. The image can includes things including but not limited to money, funding, taxes, bank, meetings with a logo of a financial institution. If you are using logo of a financial instituion to classify it as economic, make sure it is clearly visible. If it is not clearly visible, it should be classified as None’. A professional attire in itself doesn not mean economic frame.
Capacity and resources - availability of physical, human, or financial resources, and capacity of current systems. In the image, we can see things including but not limited to a geographical area, farmland, agriculture land, labour, people working in an institution, or images that convey scarcity or surplus in some way.
Morality - religious or ethical implications. In the image, we can see things including but not limited to god, death, priests, church, protests related to moral issues.
Fairness and equality - balance or distribution of rights, responsibilities, and resources. In the image, we can see things including but not limited to the fight for civil or political rights, LGBTQ, or calls to stopping discrimination.
Legality, constitutionality and jurisprudence - legal rights, freedoms, and authority of individuals, corporations, and government. In the image, we can see things including but not limited to , prisons, laws, judges in robes, courtrooms, legal documents, and prison facilities. This does not include sports contexts, such as referees or players enforcing or breaking game rules.
Policy prescription and evaluation - discussion of specific policies aimed at addressing problems. In the image, we can see things including but not limited to discussions on rule, rule making bodies, people in formal settings such as boardrooms or legislative halls - actively debating, and reviewing policy drafts or proposals. You might see official charts, graphs, or official documents. People in formal attire with no other information should not be classified as policy prescription and evaluation.
Crime and punishment - effectiveness and implications of laws and their enforcement. In the image, we can see things including but not limited to criminal activities, violence, police officers making arrests, crime scenes with investigators, courtrooms during criminal trials, prisons with detainees. This frame specifically excludes contexts involving sports, such as referees, players, or rule enforcement within games, which are not related to societal law violations or legal punishment.
Security and defense - threats to the individual, community, or nation. In the image, we can see things including but not limited to military uniforms, defense personnel, border patrol, war, soldiers, military equipment like tanks or fighter jets, border walls, or surveillance systems monitoring wide areas.
Health and safety - health care, sanitation, public safety. Images with objects like coffee, drinks, food items or activities like sports which a clear and literal message that it affects health and safety positively or negatively should be classified as health and safety, otherwise it should be classified as None’. E.g. a person drinking coffee does not mean health and safety, but a person drinking a medicine or having cigarette does. A bus does not mean health and safety, but a bus with a warning sign does. In the image, we can see things including but not limited to doctors, nurses, injury, disease, or events with environmental impact that may impact health and safety.
Quality of life - threats and opportunities for the individuals wealth, happiness, and well-being. In the image, we can see things that improves happiness or demonstrates quality of life in some form. It also includes things that demonstrate deterioration of quality of life by showing hardships of people, homelessness etc. This may also include happy children, food items that demonstrate good quality of life or people enjoying a nice meal.
Cultural identity - traditions, customs, or values of a social group in relation to a policy issue. In the image, we can see things including but not limited to concerts, cultural dance, sports, art, celebrities, artists and prominent people related to these topics. Examples, celebrities, traditional dress, sports with clear countriy specific detail e.g. jerseys/flags, culural events, cultural art etc. Otherwise, it should be classified as None’.
Public opinion - attitudes and opinions of the general public, including polling and demographics. Includes generic protests, people (non-celebrities) engaging with large crowds, riots, and strikes and including but not limited to sharing petitions and encouraging people to take political action. It will also include news broadcasts, talk shows, and interviews with people that are related to public opinion at large.
Political - considerations related to politics and politicians, including lobbying, elections, and attempts to sway voters. In the image, we can see things related to politicians, elections, voting, political campaigns. Just formal clothing does not mean political frame. If the images does not have a political person which is recognizable, it should not be classified as political. A formal attire with no political information should be classified as None’.
External regulation \& reputation - international reputation or foreign policy. In the image, we can see things including but not limited to international organizations, global discussions/meetings, foreign policy, flags from multiple countries, or delegates at a cross-country forum discussing reputation and regulation. If you use a logo of a global organization to classify it as external regulation and reputation, make sure it is clearly visible in the image. If it is not clearly visible, it should be classified as None’.
None - no frame could be identified because of lack of information in the image. This should be selected when no other frame is applicable. Example, a handshake with no other information, a logo of a company with no other information, a landscape with no other information, a person in a photo album with no other information, a person speaking with no other information about the content of the speech or persons identity, a formal event with no other information, a person in formal attire with no other information, a news logo with no news, a sports event with no additional information, simple objects like vehicle/car/pen/paper/sign-boards/objects etc with no other information etc."""
[…]
FRAMES_TASK_PROMPT = """
Given the list of frames, and the image.
<image>
Your task is to carefully analyse the image and choose the appropriate frames from the above list.
Output your answer in a json format with the format:
{"frames-list": "[<All frame names that apply from list provided above>], "reason": "<reasoning for the frames chosen>"}
Output only the json and no other text.
Listing 1: The prompt template for frame prediction using the pixtral model
[…]
TOPIC_PROMPT = "Output the topic of the article, along with a justification for the answer. The topic should be a single word or phrase. Format your output as a json entry with the field topic_justification and topic’."
ENTITY_PROMPT = """Your task is to identify the main subject or entity in the article that is central to the article. Entities are people or organizations that play a central role. If there are multiple entities being discussed, choose the one that is central to the article. If there are no central or clearly identifiable entities, choose "None".
Additionally, analyse the image and output the sentiment with which the subject is portrayed in the image. The sentiment can be "positive", "negative" or "neutral". In case of no entity, output "None. Output only the json and no other text.
Format your output as a json entry as follows:
{"entity-name": "<name of the entity>", "sentiment": "<sentiment towards the entity>", "sentiment-reason": "<reasoning for the portrayed sentiment>"}
For the given article, provide the name of the entity, sentiment towards the entity, and reasoning for the chosen sentiment."""
FRAMES = """
A list of frame names and their descriptions used in news is:
Economic - costs, benefits, or other financial implications,
Capacity and resources - availability of physical, human, or financial resources, and capacity of current systems,
Morality - religious or ethical implications,
Fairness and equality - balance or distribution of rights, responsibilities, and resources,
Legality, constitutionality and jurispudence - rights, freedoms, and authority of individuals, corporations, and government,
Policy prescription and evaluation - discussion of specific policies aimed at addressing problems,
Crime and punishment - effectiveness and implications of laws and their enforcement,
Security and defense - threats to welfare of the individual, community, or nation,
Health and safety - health care, sanitation, public safety,
Quality of life - threats and opportunities for the individuals wealth, happiness, and well-being,
Cultural identity - traditions, customs, or values of a social group in relation to a policy issue,
Public Opinion - attitudes and opinions of the general public, including polling and demographics,
Political - considerations related to politics and politicians, including lobbying, elections, and attempts to sway voters,
External regulation and reputation - international reputation or foreign policy of the U.S,
None - none of the above or any frame not covered by the above categories."""
GENERIC_FRAMING_PROMPT = f"""Framing is a way of classifying and categorizing information that allows audiences to make sense of and give meaning to the world around them (Goffman, 1974).
Entman (1993) has defined framing as "making some aspects of reality more salient in a text in order to promote a particular problem definition, causal interpretation, moral evaluation, and/or treatment recommendation for the item described".
Frames serve as metacommunicative structures that use reasoning devices such as metaphors, lexical choices, images, symbols, and actors to evoke a latent message for media users (Gamson, 1995).
A set of generic news frames with an id, name and description are: {FRAMES}.
Your task is to code articles for one of the above frames and provide justification for it. Format your output as a json entry with the fields frame_justification’, frame_id’, frame_name’.
frame_name should be one of the above listed frames. Only output one frame per article."""
GENERIC_FRAMING_MULTIPLE_PROMPT = """
Given the list of news frames, and the news article.
Your task is to carefully analyse the article and choose the appropriate frames used in the article from the above list.
Output your answer in a json format with the format:
{"frames-list": "[<All frame names that apply from list provided above>], "reason": "<reasoning for the frames chosen>"}.
Only choose the frames from the provided list of frames. If none of the frames apply, output "None" as the answer.
ISSUE_FRAMING_PROMPT = """
Entman (1993) has defined framing as "making some aspects of reality more salient in a text in order to promote a particular problem definition, causal interpretation, moral evaluation, and/or treatment recommendation for the item described".
Frames serve as metacommunicative structures that use reasoning devices such as metaphors, lexical choices, images, symbols, and actors to evoke a latent message for media users (Gamson, 1995).
There are several ways to cover a specific issue in the news. For instance, the issue of climate change can be framed as a scientific, a political, a moral, or a health issue etc. with issue-specific frames such as "Global Doom", "Local Tragedies", "Sustainable future".
Similarly, articles related to immigration can frame immigrants as a hero, a victim, or a threat with frames such as "Economic Burden", "Cultural Invasion", "Humanitarian Crisis".
Based on the topic of the article, come up with an issue-specific frame that is relevant to the topic of the article. Provide a justification for the frame.
Format your output as a json entry with the fields issue_frame_justification and issue_frame’."""
POST_PROMPT = " Output only the json and no other text. Make sure to add escape characters where necessary to make it a valid json output."
Listing 2: The prompt template for text prediction using the mistral model
Topic Name Accuracies Example
Accident 90%, 100% YONKERS, N.Y. – A 70-year-old woman was struck and killed by a car while walking on the sidewalk in Yonkers. It happened Sunday night on North Broadway. Investigators say the car then went over a retaining wall…
Crime 90%, 90% MINNEAPOLIS – The person suspected of causing a crash that killed five young women is in custody at the Hennepin County Jail. WCCO is not naming the man until he’s charged with a crime, which prosecutors say could happen as soon as Tuesday…
Culture 70%, 70% CHICAGO (CBS) – You can start your summer with a pop of color at the new Andy Warhol exhibition on the campus of the College of DuPage. With more than 200 original photographs…
Economy 100%, 100% Andrew Ross Sorkin grilled White House economic advisor Heather Boushey on Wednesday over whether the Biden administration planned for increased inflation when the president passed several spending packages. The panel was discussing….
Education 100%, 90% BALTIMORE - Westminster National Golf Course hosted 100 third graders from Westminster Elementary School on Thursday for a hands-on cross-curricular STEM-related field trip. The students learned all about golf and the science…
Entertainment 100%, 100% A new animated fantasy comedy movie that follows the adventures of a preteen Latina who wants to do her own thing while surrounded by her multigenerational Mexican American family premieres Friday on Netflix…
Finance 100%, 100% WASHINGTON, May 30 (Reuters) - The former head of Wells Fargo & Co’s (WFC.N) retail bank agreed to pay a \$3 million penalty to settle the U.S. Securities and Exchange Commission’s fraud charges for misleading investors…
Health 100%, 90% LOS ANGELES (AP)‚ Madonna has postponed her career-spanning Celebration tour due to what her manager called a “serious bacterial infection.” Manager Guy Oseary wrote on Instagram Wednesday that the singer had spent several days…
Immigration 100%, 100% A U.S. Citizenship and Immigration Services (USCIS) district office in New York City. USCIS expects to accept and approve a low number of H-1B registrations from the H-1B lottery‚ first selection round…
Legal 80%, 80% Former President Donald Trump stubbornly rejected his legal team’s efforts last year to settle the classified documents case and prevent him from being indicted by a federal grand jury, according to a bombshell report. Christopher Kise, one of Trump’s attorneys in the fall of 2022…
Lifestyle 80%, 80% The courgettes are roasting sweetly in the oven, half of them for lunch today dressed with sultanas, pine kernels and honey, the rest to serve as a salad tomorrow. This is something I also do with aubergines, red onions and sweet potatoes. There are so many vegetables to roast right now…
Politics 90%, 100% Rep. Matt Gaetz (R-Fla.) criticized House Republicans’ recent effort to impeach President Joe Biden saying it was not done in a “legitimate” or “serious” way, a video obtained by NBC News shows, raising questions of whether he will support Rep. Jim Jordan (R-Ohio), who is a leading candidate to become House speaker…
Safety 90%, 50% The Sacramento region has some of the highest numbers of fatal traffic collisions in the state. Sacramento police say that last year, more than 50 people died on city streets. Now, as part of National Passenger Safety Week, there’s an effort to reduce fatal collisions…
Social Issues 70%, 70% REDWOOD CITY, The San Mateo County Board of Supervisors is continuing to explore ways to provide more housing for farmworkers in the county, nearly four months after a mass shooting in Half Moon Bay exposed an urgent need for more living options for agricultural workers with low income…
Technology 70%, 100% America’s spending on artifical intelligence in public safety is projected to increase from $9.3billionin2022to$71 billion by 2030, according to a new analysis by the Insight Partners research firm. The projected seven-year boom is expected to be fueled by global and domestic terrorism…
War 100%, 100% The Ukrainian air-assault force‚Äôs 25th Brigade just got a lot heavier. Photos that appeared online this week confirm the brigade has re-equipped with German-made Marder infantry fighting vehicles. The 25th is the second Ukrainian air-assault brigade, after the 82nd, to get 31-ton Marders from German stocks…
Weather 100%, 100% Tropical Storm Nigel is expected to become a hurricane as soon as Monday, the National Hurricane Center said Sunday, and could be the latest tropical storm in the Atlantic this season to rapidly intensify to major hurricane status…
Table 7: The 19 topics used in our analysis. Accuracies represent the average acceptability judgments of two annotators (two of the paper authors) over a set of 10 predicted examples for each topic.